### v2

# KAN or MLP? Point Cloud Shows the Way Forward

Abstract
Multi-Layer Perceptron (MLP) have become one of the fundamental architectural component in point cloud analysis due to its effective feature learning mechanism. However, when processing complex geometric structures in point clouds, MLP’ fixed activation functions struggle to efficiently capture local geometric features, while suffering from poor parameter efficiency and high model redundancy. In this paper, we propose PointKAN, which applies Kolmogorov-Arnold Network (KAN) to point cloud analysis tasks to investigate their efficacy in hierarchical feature representation. First, we introduce a Geometric Affine Module (GAM) to transform local features, improving the model’s robustness to geometric variations. Next, in the Local Feature Processing (LFP), a parallel structure extracts both group-level features and global context, providing a rich representation of both fine details and overall structure. Finally, these features are combined and processed in the Global Feature Processing (GFP). By repeating these operations, the receptive field gradually expands, enabling the model to capture complete geometric information of the point cloud. To overcome the high parameter counts and computational inefficiency of standard KAN, we develop Efficient-KAN in the PointKAN-elite variant, which significantly reduces parameters while maintaining accuracy. Experimental results demonstrate that PointKAN outperforms PointMLP on benchmark datasets such as ModelNet40, ScanObjectNN, and ShapeNetPart, with particularly strong performance in Few-shot Learning task. Additionally, PointKAN achieves substantial reductions in parameter counts and computational complexity (FLOPs). This work highlights the potential of KAN-based architectures in 3D vision and opens new avenues for research in point cloud understanding.
Figure 1: PointKAN (including PointKAN-elite) comprehensively outperforms counterparts. (a) Surpasses MLP and CNN baselines across datasets; (b) In Few-shot Learning, exceeds Point-BERT’s pre-trained performance when trained from scratch; (c)-(d) Substantially reduces parameters versus PointMLP, with PointKAN-elite significantly curbing FLOPs growth as Patch Sequence Length increases.
1 Introduction
Point cloud analysis is vital in computer vision and 3D processing, with applications spanning autonomous driving  (Li, Luo, and Yang 2023) and robotics (Yan et al. 2020; Rusu and Cousins 2011). Unlike structured 2D images, point clouds pose challenges due to their irregularity and sparsity. Recent deep learning advances (Zhang et al. 2022; Pang et al. 2022; He et al. 2024; Liang et al. 2024; Han et al. 2024) build upon pioneering MLP-based architectures like PointNet/PointNet++  (Qi et al. 2017a, b). Subsequent research, including PointNeXt  (Qian et al. 2022), DGCNN  (Wang et al. 2019), and Point Transformer  (Zhao et al. 2021), employs CNN, GCN, and attention to capture local geometric features.
Notably, PointMLP (Ma et al. 2022) utilizes a pure MLP architecture for point cloud analysis. Its core design philosophy eschews complex convolutions or graph operations, instead relying on stacked residual MLP blocks for feature learning. While achieving competitive performance in classification and segmentation, PointMLP inherits inherent MLP limitations. Rooted in the Universal Approximation Theorem (Cybenko 1989), MLPs employ linear weights and fixed activations. Consequently, capturing intricate geometric relationships necessitates scaling network dimensions, leading to parameter explosion and scaling challenges.
In contrast, the recently proposed Kolmogorov-Arnold Network (KAN)  (Liu et al. 2024) offers a promising MLP alternative, gaining significant attention for its accuracy and parameter efficiency. KAN leverages the Kolmogorov-Arnold Representation Theorem (KART)  (Arnold 2009), which states that any multivariate continuous function can be decomposed into univariate functions. KAN replaces fixed activation functions with learnable univariate functions (e.g., B-splines), directly modeling non-linearities and reducing reliance on redundant linear layers. This makes KAN particularly suitable for capturing hierarchical geometric features in irregular data like point clouds.
Building on this foundation, we propose PointKAN, a KAN-based architecture for point cloud analysis. Point clouds are mapped to high-dimensional space and processed by a Geometric Affine Module (GAM). GAM utilizes Group-Norm for point grouping, normalization, and affine transformation, alongside S-Pool (employing Softmax-Pooling) to aggregate group-wise global information. These parallel outputs feed into KAN layers to extract complementary local and global features. Local features undergo max-pooling and are fused with global information, followed by a Residual Point (ResP) Block for deep feature aggregation. To address KAN’s computational inefficiency and parameter growth, PointKAN-elite integrates Efficient-KAN, inspired by Group-Rational KANs  (Yang and Wang 2024). Trained from scratch, PointKAN achieves excellent performance across point cloud datasets (Fig 1(a)), surpassing even pre-trained Point-BERT  (Yu et al. 2022) in few-shot learning (Fig 1(b)). It also demonstrates significant reductions in parameters and FLOPs (Fig 1(c-d)), highlighting KAN’s potential for 3D point cloud processing.
In summary, this work makes the following contributions.
  • We introduce the Kolmogorov-Arnold Network (KAN) for point cloud analysis, named PointKAN, which has a stronger ability to learn complex geometric features.
  • We propose PointKAN-elite, an elite version of PointKAN, which significantly reduces both parameter count and computational cost while maintaining comparable experimental results to the original PointKAN.
  • During hierarchical point cloud processing, we propose S-Pool to aggregate features within groups, avoiding information loss while providing global context for each group of points.
2 Related Work
2.1 A Review of PointMLP
Traditional point cloud analysis methods (e.g., graph convolution  (Wang et al. 2019) or attention mechanisms  (Zhao et al. 2021)) involve complex local geometric operations, resulting in high computational costs and limited generalization capability.
PointMLP  (Ma et al. 2022) simplifies network architecture to enhance efficiency and performance. It employs a multi-stage design where Farthest Point Sampling (FPS) and K-Nearest Neighbors (KNN) progressively reduce point density at each stage. Local features are extracted via Residual MLP Blocks—comprising linear layers, normalization, and activations with residual connections to mitigate gradient vanishing. Stacking these stages forms a deep hierarchical network for point cloud analysis. The core operation is formulated as:
$g_{i}=\Phi_{pos}(\mathcal{A}(\Phi_{pre}(f_{i,j}),|j=1,\cdots,K)),$ (1)
where $\Phi_{pre}$ and $\Phi_{pos}$ are residual MLP blocks, $\mathcal{A}$ denotes max-pooling and $f_{i,j}$ represents the feature of the $j$-th point in the $i$-th group, with each group encompassing $K$ points. Despite its simple architecture, PointMLP (Ma et al. 2022) delivers excellent performance on benchmarks like ModelNet40 (Wu et al. 2015) and ScanObjectNN (Uy et al. 2019).
2.2 Kolmogorov-Arnold Network
Kolmogorov-Arnold Networks (KAN)  (Liu et al. 2024) represent a novel architecture grounded in the Kolmogorov-Arnold Representation Theorem (KART)  (Arnold 2009). Unlike MLP, KAN replaces fixed activation functions with learnable univariate functions. KART establishes that any multivariate continuous function decomposes into finite univariate functions. KAN implements this using B-spline bases and compositional operations, providing a flexible and interpretable framework for high-dimensional function approximation.
KAN demonstrates growing efficacy across domains including computer vision  (Ferdaus et al. 2024; Ma et al. 2025; Moradi et al. 2024; Mohan, Wang, and Zhu 2024) and medical imaging  (Yang et al. 2025; Agrawal et al. 2024). In vision, KAN-augmented convolutional layers replace linear transformations with learnable per-pixel nonlinear activations  (Bodner et al. 2024), enhancing parameter richness and expressiveness for image recognition. Medical implementations integrate KAN layers into frameworks like U-Net  (Ronneberger, Fischer, and Brox 2015), improving segmentation accuracy and interpretability  (Wu et al. 2024; Li et al. 2024; Tang, Chen, and Shu 2024). For point clouds, PointNet-KAN  (Kashefi 2024) shows competitive classification and segmentation results versus PointNet  (Qi et al. 2017a). However, PointNet’s inherent limitations constrain this benchmark, and KAN’s potential for complex point cloud tasks remains largely unexplored. Effective integration of KAN with advanced point cloud architectures thus presents a critical research direction.
3 Method
We propose leveraging KAN for joint local-global feature learning in point clouds. To achieve this, we first review the Kolmogorov-Arnold Representation Theorem (KART) and KAN fundamentals(section 3.1), then detail our model’s key designs(section 3.2) including Efficient-KAN(section 3.3) in PointKAN-elite.
3.1 Preliminaries: KART and KAN
The Kolmogorov-Arnold Representation Theorem (KART), established by Andrey Kolmogorov and Vladimir Arnold in the 1950s, asserts that any continuous multivariate function $f:[0,1]^{d}\rightarrow\mathbb{R}$,
$f(x_{1},x_{2},\ldots,x_{d})=\sum_{q=1}^{2d+1}\phi_{q}\left(\sum_{p=1}^{d}\psi_{q,p}(x_{p})\right),$ (2)
where $\psi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$ are univariate continuous functions. Although eq. 2 theoretically enables high-dimensional function representation via two univariate layers, these functions may be potentially highly complex (even fractal) and thus unlearnable by neural networks.
KAN’s innovation overcomes KART’s two-layer function limitation by implementing a multi-layer network trained via backpropagation. A continuous and smooth univariate spline function $spline(x)$ is generated by the linear combination of B-spline base functions ($B_{i}(x)$), as shown in eq. 3. The coefficients $c_{i}$ therein are learnable. The activation function $\phi(x)$ is the sum of $silu(x)$ and $spline(x)$, where $silu(x)$ acts similar to a residual connection, as shown in eq. 4.
$spline(x)=\sum_{i}c_{i}B_{i}(x),$ (3)
$\phi(x)=silu(x)+spline(x),$ (4)
by replacing traditional fixed activation functions with this approach, the model’s expressiveness and efficiency are enhanced while maintaining its interpretability.
Figure 2: Illustration of PointKAN. (a) Each of the four identical stages processes local features through a Geometric Affine Module (GAM), extracts grouped features via Local Feature Processing (LFP), supplements global information after aggregation, and obtains overall features through Global Feature Processing (GFP). This multi-stage repetition progressively enlarges the receptive field for comprehensive geometric capture. (b) Group-Norm comprises normalization, affine transformation, and feature concatenation; (c-d) present KAN Block and Layer details; (e) The Residual Point (ResP) Block integrates MLP with Batch Normalization and ReLU activation.
3.2 PointKAN
Despite KAN’s superior high-dimensional approximation capabilities and parameter efficiency, its adaptation to 3D point clouds remains challenging. The discrete spline nodes in KAN’s activation functions struggle to comprehensively capture local geometric features, while per-dimension parameter storage incurs significant memory overhead for large-scale networks. Additionally, B-spline computations exhibit suboptimal hardware acceleration, leading to inference latency. To overcome these limitations, we propose PointKAN—featuring a geometric affine module and parallel local feature extraction structure—alongside its lightweight variant PointKAN-elite, which jointly reduce memory footprint and accelerate training and inference speed. The complete pipeline is illustrated in fig. 2.
Geometric Affine Module
This module integrates Group-Norm and S-Pool for enhanced local feature extraction. Given input point cloud data $P\in\mathbb{R}^{N\times 3}$, we first apply Farthest Point Sampling (FPS) to select $G$ center points $P_{C}\in\mathbb{R}^{G\times 3}$, then construct local groups $\{x^{i}_{p}\in\mathbb{R}^{K\times 3}\}_{i=1}^{G}$ around each center using K-Nearest Neighbors (KNN). A lightweight PointNet (Qi et al. 2017a) processes each group to generate features $\{f^{i}_{j}\}_{j=1}^{K}\in\mathbb{R}^{K\times d}$ (where $d$ denotes feature dimension) and center features $\{f^{i}\}_{i=1}^{G}\in\mathbb{R}^{G\times d}$, which collectively form the Group-Norm input.
To enhance model stability and robustness, we design the Group-Norm module performing intra-group feature normalization, learnable affine transformation, and central-point feature propagation. Features $\{f^{i}_{j}\}_{j=1}^{K}$ are normalized:
$\hat{\{f^{i}_{j}\}}=\frac{\{f^{i}_{j}\}-f^{i}}{\sqrt{Var(\{f^{i}_{j}\}-f^{i})+\epsilon}},$ (5)
yielding $\mathbb{R}^{K\times d}$ features. We then apply an affine transformation $\tilde{f^{i}_{j}}=\alpha\odot\hat{f^{i}_{j}}+\beta$ with learnable parameters $\alpha,\beta\in\mathbb{R}^{2d}$ ($\odot$: Hadamard product), followed by concatenation $\tilde{f^{i}_{j}}\oplus f^{i}$ ($\oplus$ : feature-dimension concatenation) to propagate central-point features, outputting $\mathbb{R}^{K\times 2d}$ features. Here $\epsilon=1e^{-5}$ ensures numerical stability, where the affine transformation captures rigid feature transformations while preserving local characteristics.
To address the lack of global context from mutually independent group features when directly feeding Group-Norm outputs to KAN, we introduce S-Pool in the GAM. This aggregates group-wise features, with S-Pool and Group-Norm outputs processed in parallel to supplement global information. Unlike max-pooling which maintains permutation invariance but incurs information loss, our softmax-inspired S-Pool maximizes intra-group feature preservation as defined in eq. 6,
$\tilde{f^{i}}={\textstyle\sum_{j}^{K}}\frac{\exp{\tilde{f^{i}_{j}}}}{\sum_{k}^{K}\exp{\tilde{f^{i}_{k}}}}\cdot\tilde{f^{i}_{j}},$ (6)
where $\tilde{f^{i}}$ denotes the updated feature of the group center. This S-Pool method proportionally integrates multi-dimensional features of each point within a group into the final center feature, achieving a mapping from $\{\tilde{f^{i}_{j}}\}_{j=1,\cdots,K}\in\mathbb{R}^{G\times K\times 2d}\mapsto\tilde{f^{i}}\in\mathbb{R}^{G\times 2d}$.
Local Feature Processing
In the Local Feature Processing (LFP), the two outputs of the GAM are processed in parallel using the KAN block. The KAN block, containing the KAN layer, effective depth-wise convolution (Chollet 2017) layer (DwConv), and residual connection, adds DwConv after the KAN layer to help KAN learn rich feature representations from high-dimensional channel information. Let $\mathcal{F}(\cdot)$ represent the processing of the KAN block, which can be denoted as:
$\mathcal{F}({f_{in}})={f_{in}}+DwConv(\Phi({f_{in}})),$ (7)
where ${f_{in}}$ denotes the input point cloud features, $\Phi$ represents the KAN layer. The shared LFP module is designed to learn local features from grouped point clouds. After max-pooling, it is added to the global information learned from the central point clouds of each group. This makes the aggregated feature information outputted by each group of point clouds more abundant.
Global Feature Processing
The Global Feature Processing (GFP) module utilizes Residual Point (ResP) Blocks to extract deep aggregated features, leveraging MLP’s inherent capacity for cross-channel fusion through fully-connected layers—a capability where KAN exhibits limitations despite their strength in capturing high-frequency details via univariate function decomposition. The ResP Block’s structure $x+MLP(x)$, with MLP comprising two sequential stacks of linear layers, normalization, and activations, provides critical advantages: its permutation invariance satisfies point cloud processing requirements, residual connections enable deep feature extraction, and purely feed-forward operations ensure computational efficiency for massive data. These properties facilitate efficient stacking of multiple ResP Blocks within GFP, as validated by ablation studies (see supplementary materials).
In general, instead of using complex local geometric extraction structures, one stage of PointKAN consists of three components: the Geometric Affine Module, the Local Feature Processing, and the Global Feature Processing. We construct a deep network for hierarchical point-cloud processing through four recursively repeated stages.
3.3 Efficient-KAN
KAN’s B-spline activation functions $\phi(x)$ exhibit computational inefficiency due to recursive computation and poor GPU parallelization. Each input-output pair requires dedicated parameters and base functions, causing exponential parameter growth with hidden layer width. To mitigate these limitations, we introduce Efficient-KAN in PointKAN-elite as a scalable alternative.
In comparison to KAN, Efficient-KAN utilize rational functions as the base functions within KAN, replacing the B-spline functions. Specifically, each activation function $\phi(x)$ is computed using rational polynomials $P(x)$ and $Q(x)$ as defined in eq. 8, where the degrees of $P(x)$ and $Q(x)$ are denoted as $m$ and $n$, respectively.
$\phi(x)=wF(x)=w\frac{P(x)}{Q(x)}=w\frac{a_{0}+a_{1}x+\cdots+a_{m}x^{m}}{\sqrt{1+(b_{1}x+\cdots+b_{n}x^{n})^{2}}},$ (8)
where $\{a_{i}\}_{i=1,\cdots,m}$ and $\{b_{j}\}_{j=1,\cdots,n}$ are coefficients of the rational functions, and $w$ is the scaling factor. These parameters are trained via backpropagation, where rational functions implemented with CUDA demonstrate lower computational complexity than B-spline functions and enhanced parallelism. We specifically implement two corresponding optimizations. First, we employed Horner’s method (Horner 1815) for polynomial evaluation to further reduce computational overhead. Horner’s method is described as follows:
$a_{0}+a_{1}x+\cdots+a_{m}x^{m}=a_{0}+x\left(a_{1}+x\left(a_{2}+x(\cdots)\right)\right),$ (9)
which evaluates a polynomial of degree $m$ using only $m$ multiplications and $m$ additions. Subsequently, the explicit gradients of $F(x)$ with respect to $\frac{\partial F}{\partial a_{m}}$, $\frac{\partial F}{\partial b_{n}}$ and $\frac{\partial F}{\partial x}$ can be computed respectively:
$\quad\frac{\partial F}{\partial a_{m}}=\frac{x^{m}}{Q(x)},\quad\frac{\partial F}{\partial b_{n}}=-x^{n}\frac{G(x)}{Q^{3}(x)}P(x),$ (10)
$\quad\frac{\partial F}{\partial x}=\frac{\partial P(x)}{\partial x}\frac{1}{Q(x)}-\frac{\partial Q(x)}{\partial x}\frac{P(x)}{Q^{2}(x)}$
which $G(x)=b_{1}x+\cdots+b_{n}x^{n}$, $\frac{\partial P(x)}{\partial x}=a_{1}+2a_{2}x+\cdots+ma_{m}x^{m-1}$ and $\frac{\partial Q(x)}{\partial x}=\frac{G(x)}{Q(x)}\left(b_{1}+2b_{2}x+\cdots+nb_{n}x^{n-1}\right)$. All factors in the result can be further accelerated computationally using Horner’s method.
Another significant change in Efficient-KAN is the grouping of input channels, which allows parameter sharing within groups to reduce the number of parameters and computational load. As illustrated in fig. 3, MLP utilizes fixed activation functions, while KAN assigns an activation function to each input-output pair. In contrast, Efficient-KAN integrates the approaches of the former two by sharing activation functions within groups to minimize the number of parameters.  Table 1 presents a quantitative comparison of the parameter counts among the single-layer MLP, the Vanilla KAN, and Efficient-KAN, in which $d_{in}$ and $d_{out}$ represent the input and output dimensions, respectively; $k$ denotes the spline order; $G$ indicates the grid size; $m$ and $n$ correspond to the polynomial degrees of rational functions; and $g$ is the number of groups. The parameter count of Efficient-KAN introduces only a constant term compared to MLP, whereas original KAN’s parameter scale grows roughly as (G+k+2) times that of MLP.
Figure 3: Comparing among MLP, Vanilla KAN, and Efficient-KAN. MLP: each output channel uses the same activation function. Vanilla KAN: each input-output pair has distinct activation functions. Efficient-KAN: input-output pairs within the same group employ the same activation function.
Table 1: The parameter comparison among single-layer MLP, Vanilla KAN, and Efficient-KAN.
Model Parameter count
MLP $d_{\text{in}}\times d_{\text{out}}+d_{\text{out}}$
Vanilla KAN $d_{\text{in}}\times d_{\text{out}}\times(G+k+2)+d_{\text{out}}$
Efficient-KAN (Ours) $d_{\text{in}}\times d_{\text{out}}+d_{\text{out}}+(n+m\times g)$
4 Experiments
In this section, we first introduce detailed settings of experiments. Subsequently, PointKAN is compared and evaluated against various existing methods, especially PointMLP, on multiple downstream tasks, including object classification, part segmentation, and few-shot learning. Finally, we present results of ablation studies.
4.1 Implementation Details
Experiments are conducted on three benchmarks using a single NVIDIA RTX 4090: ModelNet40 (Wu et al. 2015) and ScanObjectNN (Uy et al. 2019) for classification, and ShapeNetPart (Yi et al. 2016) for part segmentation, with hyperparameters aligned to PointMLP (Ma et al. 2022) for fair comparison. The PointKAN architecture employs four identical stages per task, each containing a GAM, LFP, and GFP module, using a batch size of 18. For classification, SGD optimizes inputs of 1,024 points grouped into 24 per partition; for segmentation, Adam with cosine decay handles 2,048 points (including normals) grouped into 32 per partition. The PointKAN-elite variant increases the batch size to 32 while halving learning rates, retaining other configurations. All ablation studies default to PointKAN and are evaluated primarily on ModelNet40.
Table 2: Classification results on ModelNet40 and ScanObjectNN. The inference model parameters Param.(M), FLOPs(G), and overall accuracy (%) are reported. We compare the KAN-based model with methods based on CNN, GCN, Transformer, Mamba, and MLP.
Method Param.(M) $\downarrow$ FLOPs(G) $\downarrow$ ModelNet40 ScanObjectNN
1k P $\uparrow$ OBJ_BG $\uparrow$ OBJ_ONLY $\uparrow$ PB_T50_RS $\uparrow$
Based on CNN or GCN
DGCNN (Wang et al. 2019) 1.8 2.4 92.9 82.8 86.2 78.1
PointCNN (Li et al. 2018) 0.6 - 92.2 86.1 85.5 78.5
GBNet (Qiu, Anwar, and Barnes 2021b) 8.8 - 93.8 - - 81.0
MVTN (Hamdi, Giancola, and Ghanem 2021) 11.2 43.7 93.8 92.6 92.3 82.8
Based on Transformer or Mamba
Transformer (Vaswani et al. 2017) 22.1 4.8 91.4 79.86 80.55 77.24
PCT (Guo et al. 2021) 2.9 2.3 93.2 - - -
PointMamba (Liang et al. 2024) 12.3 3.6 - 88.30 87.78 82.48
PCM (Zhang et al. 2024) 34.2 45.0 93.4$\pm$0.2 - - 88.1$\pm$0.3
Based on MLP
PointNet (Qi et al. 2017a) 3.5 0.5 89.2 73.3 79.2 68.0
PointNet++ (Qi et al. 2017b) 1.5 1.7 90.7 82.3 84.3 77.9
DRNet (Qiu, Anwar, and Barnes 2021a) - - 93.1 - - 80.3
SimpleView (Goyal et al. 2021) - - 93.9 - - 80.5$\pm$0.3
PRA-Net (Cheng et al. 2021) 2.3 93.7 - - 81.0
PointMLP (Ma et al. 2022) 12.6 31.4 93.3 - - 85.0$\pm$0.3
Based on KAN
PointNet-KAN (Kashefi 2024) - - 90.5 - - 60.2
PointKAN (Ours) 10.0 9.1 93.7 87.1 86.4 83.5
PointKAN-elite (Ours) 3.1 2.3 93.3 84.3 85.3 84.1
Table 3: Few-shot classification on ModelNet40. Overall accuracy (%) without voting and pre-training strategies is reported.
Method 5-way 10-way
10-shot $\uparrow$ 20-shot $\uparrow$ 10-shot $\uparrow$ 20-shot $\uparrow$
Based on CNN
DGCNN (Wang et al. 2019) 31.6 $\pm$ 2.8 40.8 $\pm$ 4.6 19.9 $\pm$ 2.1 16.9 $\pm$ 1.5
DGCNN+OcCo (Wang et al. 2021) 90.6 $\pm$ 2.8 92.5 $\pm$ 1.9 82.9 $\pm$ 1.3 86.5 $\pm$ 2.2
DGCNN-CrossPoint (Afham et al. 2022) 92.5 $\pm$ 3.0 94.9 $\pm$ 2.1 83.6 $\pm$ 5.3 87.9 $\pm$ 4.2
Based on Transformer or Mamba
Transformer (Vaswani et al. 2017) 87.8 $\pm$ 5.2 93.3 $\pm$ 4.3 84.6 $\pm$ 5.5 89.4 $\pm$ 6.3
MAMBA3D  (Han et al. 2024) 92.6 $\pm$ 3.7 96.9 $\pm$ 2.4 88.1 $\pm$ 5.3 93.1 $\pm$ 3.6
Point-BERT (Yu et al. 2022) 94.6 $\pm$ 3.1 96.3 $\pm$ 2.7 91.0 $\pm$ 5.4 92.7 $\pm$ 5.1
MaskPoint (Liu, Cai, and Lee 2022) 95.0 $\pm$ 3.7 97.2 $\pm$ 1.7 91.4 $\pm$ 4.0 93.4 $\pm$ 3.5
Based on MLP
PointNet (Qi et al. 2017a) 52.0 $\pm$ 3.8 57.8 $\pm$ 4.9 46.6 $\pm$ 4.3 35.2 $\pm$ 4.8
PointNet-CrossPoint (Afham et al. 2022) 90.9 $\pm$ 1.9 93.5 $\pm$ 4.4 86.4 $\pm$ 4.7 90.2 $\pm$ 2.2
PointMLP (Ma et al. 2022) 90.5 $\pm$ 4.2 94.2 $\pm$ 4.1 84.1 $\pm$ 5.6 91.5 $\pm$ 5.1
Based on KAN
PointKAN(Ours) 95.9 $\pm$ 3.1 97.9 $\pm$ 2.0 90.5 $\pm$ 4.9 94.0 $\pm$ 3.5
PointKAN-elite(Ours) 95.2 $\pm$ 3.1 96.3 $\pm$ 3.0 88.0 $\pm$ 5.1 93.5 $\pm$ 4.0
4.2 Experimental Results on Downstream Tasks
PointKAN (including PointKAN-elite) demonstrates inherent advantages over PointMLP (Ma et al. 2022) in multi-dimensional benchmarking across downstream tasks. Our models are trained from scratch for each specific task, with no implementation of pre-training methodologies or ensemble voting mechanisms in any experimental configuration.
Synthetic Object Classification on ModelNet40
The ModelNet40 (Wu et al. 2015) dataset constitutes a comprehensive and clean collection of 3D CAD models, encompassing 12,311 human-annotated models spanning 40 distinct categories. The dataset is partitioned into 9,843 models for training and 2,468 models for testing. On ModelNet40, PointKAN achieves 93.7% OA (+0.4% vs PointMLP) with 21% fewer parameters (10M vs 12.6M) and 71% lower FLOPs (9.1G vs 31.4G). Furthermore, PointKAN-elite achieves remarkable parameter efficiency (3.1M parameters) and computational efficiency (2.3G FLOPs), while maintaining competitive performance with 90.7% mAcc and 93.3% OA. Results shown in table 2.
Real-world Object Classification on ScanObjectNN
The ScanObjectNN dataset (Uy et al. 2019), derived from 3D scans of real-world indoor scenes, comprises 15,000 point cloud object samples spanning 15 daily object categories (e.g., chairs, tables, monitors) and encompassing 2,902 distinct object instances. On ScanObjectNN, 15 categories), PointKAN achieves 87.1% (OBJ_BG) and 86.4% (OBJ_ONLY) OA in 200 epochs, outperforming scratch-trained baselines. For PB_T50_RS, it attains 83.5% OA (vs PointMLP’s 85.0%), while PointKAN-elite reaches 84.1% with reduced computation. Results shown in table 2.
Few-shot Learning
Following the $n$-way $m$-shot protocol ($n{\in}\{5,10\}$, $m{\in}\{10,20\}$) on ModelNet40, PointKAN achieves superior few-shot accuracy versus CNN and Transformer baselines (see table 3), outperforming MAMBA3D (Han et al. 2024) and Point-BERT (Yu et al. 2022) in most settings. Excluding the 10-way 10-shot case (marginally below MaskPoint (Liu, Cai, and Lee 2022)), it shows consistent OA gains over PointMLP (+5.4%, +3.7%, +6.4%, and +2.5%). Notably, its scratch-trained performance exceeds Point-BERT’s pre-trained results, demonstrating strong generalization.
Part Segmentation on ShapeNetPart
ShapeNetPart (Yi et al. 2016) to predict more fine-grained category labels for each point within the samples. The ShapeNetPart dataset comprises 16,881 3D instances spanning 16 object categories with a total of 50 part labels. We evaluate part segmentation on ShapeNetPart. As shown in  table 4, PointKAN achieves competitive mIoU versus PointMLP with fewer parameters (13.8M vs 16.0M) and FLOPs (3.8G vs 5.8G). PointKAN-elite further improves efficiency (6.2M vs. 16.0M, 1.4G vs. 5.8G). We also provide a comparative visualization between the segmentation ground truths and predicted outcomes in  fig. 4, demonstrating KAN’s potential. Additional part segmentation visualization results are provided in the supplementary materials.
Table 4: Part segmentation results on ShapeNetPart. The class mIoU (Cls. mIoU) and the instance mIoU (Inst. mIoU) are reported.
Method
Cls.
mIoU
$\uparrow$
Inst.
mIoU
$\uparrow$
Param.(M)$\downarrow$ FLOPs(G) $\downarrow$
Based on CNN or GCN
PCNN (Atzmon, Maron, and Lipman 2018) 81.8 85.1 5.4 -
DGCNN (Wang et al. 2019) 82.3 85.2 1.3 12.4
PointCNN (Li et al. 2018) 84.6 86.1 - -
RS-CNN (Liu et al. 2019) 84.0 86.2 - -
SpiderCNN (Xu et al. 2018) 82.4 85.3 - -
Based on MLP
PointNet (Qi et al. 2017a) 80.4 83.7 3.6 4.9
PointNet++ (Qi et al. 2017b) 81.9 85.1 1.0 4.9
PointMLP (Ma et al. 2022) 84.6 86.1 16.0 5.8
Based on KAN
PointKAN(Ours) 83.9 85.6 13.8 3.8
PointKAN-elite(Ours) 84.3 86.0 6.2 1.4
Figure 4: Part segmentation results on ShapeNetPart. Top line is ground truth and bottom line is our prediction.
4.3 Ablation Study
We perform comprehensive ablation studies on the architecture and KAN parameters. All results are obtained from models trained from scratch on ModelNet40, with additional analyses provided in supplementary materials.
Architecture Ablation
Architecture ablation studies ( table 5) demonstrate that integrating S-Pool in GAM elevates overall accuracy by 0.7% through compensating global information loss during local grouping via parallel processing with Group-Norm; conversely, removing LFP reduces OA by 1.0%, validating KAN’s critical role in local feature extraction for unordered point clouds, while omitting the Residual Point Block in GFP decreases accuracy by 0.5%, confirming its necessity for global feature aggregation. The full component integration achieves peak 93.7% OA.
Ablation of Built-in Parameters in KAN
KAN critically determines model performance. Firstly, we analyze two parameters, $k$ (the order of the B-spline function) and $G$ (the number of grids) in  fig. 5. Higher $k$ enhances curve continuity while $G$ controls knot interval density, jointly affecting flexibility. However, excessive values cause overfitting and computational complexity.
In addition, guided by KAN’s generalization principles  (Liu et al. 2024), we balance network scale: insufficient capacity impairs feature learning, while oversized networks harm generalization. Through architectural experiments ( table 6), a three-layer KAN with $k=3$, $G=5$, and hidden dimension $d_{hid}=d_{in}/2$ achieves optimal accuracy.
5 Conclusion
We propose PointKAN, an efficient KAN-based architecture for point cloud analysis. PointKAN outperforms PointMLP across multiple tasks, demonstrating superior local feature extraction. The lightweight PointKAN-elite variant further maintains accuracy while substantially reducing parameters and FLOPs. We anticipate PointKAN will advance KAN applications in point cloud analysis and leverage its unique advantages over MLP.
Table 5: Component ablation studies on ModelNet40 test set. The mean accuracy (%) and overall accuracy (%) are reported.
S-Pool LFP GFP mAcc(%) OA(%)
90.8 93.0
89.9 92.7
90.7 93.2
91.4 93.7
Figure 5: Ablation of Built-in Parameters $k$ and $G$ in KANs.The mean accuracy (%) and overall accuracy (%) are reported.
Table 6: Ablation on KANs’ depth and hidden layer width. $d_{in}$ denotes the input feature dimension. The mean accuracy (%) and overall accuracy (%) are reported.
Depth Width mAcc(%) OA(%)
3 layers $d_{in}\times{1/3}$ 90.6 93.0
$d_{in}\times{1/2}$ 91.4 93.7
$d_{in}\times{2/3}$ 90.5 92.8
$d_{in}$ 90.4 93.5
4 layers $d_{in}\times{1/3}$ 90.1 92.9
$d_{in}\times{1/2}$ 91.0 93.2